A parallel-plate capacitor has a plate area of 2.5

a= 7*10^-8

b = 1.4*10^-6
Answer #1

a)

Voltage across the capacitor

V=Ed =(3*106)(2*10-3)

V=6000 Volts

Capacitance

C=eoA/d =(8.85*10-12)(2.5*10-3)/(2*10-3)

C=1.11*10-11F

Maximum charge

Qmax=CV =(1.11*10-11)*6000

Qmax=6.64*10-8C =7*10-8 C (approx)

b)

Voltage across the capacitor

V=Ed =(24*106)(2*10-3)

V=48000 Volts

Dielectric constant k=2.56

Capacitance

Cp=KC =2.84*10-11F

Maximum charge

Qmax=CV =(2.84*10-11)*48000

Qmax=1.36*10-6C =1.4*10-6 C (approx)